OpenStudy (anonymous):

use the given information to find f ' (2) g(2)= 2 g ' (2)= -4 h(2)= -1 h ' (2) = 4 f(x)= g(x)/h(x)

OpenStudy (anonymous):

i know you use the chain rule h ' (x)= f ' [g(x)] * g ' (x) im just a little confused on how

OpenStudy (amistre64):

f is a quotient, what is the quotient rule?

OpenStudy (amistre64):

this isnt a chain rule perse

OpenStudy (anonymous):

the quotient rule is v/u= vu' - uv'/(v)^2 but our teacher shows us problem like this using the other rule i put on my 1st reply not the quotient rule

OpenStudy (amistre64):

this is a quotient, use the quotient rule

OpenStudy (amistre64):

f = g/h f' = (gh' - g'h)/h^2 they give you g,h,g',h' to fill in

OpenStudy (anonymous):

thats what i thought but when a kid ask about using the quotient rule in class he told us not quit thats why i was confused

OpenStudy (amistre64):

if f=g(h) we use a chain rule if f = gh we use a product rule if f = g/h we use a quotient rule if f = g+h we use whatever rules apply to g and h

OpenStudy (anonymous):

but thank you ill try your way

OpenStudy (anonymous):

yeah web assign told me it was wrong maybe i did something wrong who knows

OpenStudy (amistre64):

2(4) - (-1)(-4) ------------ (-1)^2 so 8-4 is?

OpenStudy (amistre64):

well, 4-8

OpenStudy (anonymous):

4 but web assign said that was wrong

OpenStudy (amistre64):

i got my top quotient rule backwards ... hg' - h'g

OpenStudy (anonymous):

ok yeah i did 8-4 not 4-8

OpenStudy (amistre64):

i usually define it as but but t/b = bt'-b't/b^2

OpenStudy (anonymous):

thank you

OpenStudy (amistre64):

youre welcome